OpenStudy (lynfran):

Help!! Please.. Prove that (1-tan^2(pi/4 - theta)) divided by (1 + tan^2(pi/4 - theta)) = 2sin(theta)cos(theta)

OpenStudy (zehanz):

It certainly looks horrible! : \(\dfrac{1-\tan^2 (\frac{\pi}{4}-\theta)}{1+\tan^2 (\frac{\pi}{4}-\theta)}=2\sin \theta \cos \theta\). I will do the first step, then you can try the rest. Remember, \(\tan \theta=\dfrac{\sin \theta}{\cos \theta }\), so the LHS of the identity becomes: \(\dfrac{1-\dfrac{\sin^2 (\frac{\pi}{4}-\theta)}{\cos^2(\frac{\pi}{4}-\theta)}}{1+\dfrac{\sin^2 (\frac{\pi}{4}-\theta)}{\cos^2(\frac{\pi}{4}-\theta)}} \).

OpenStudy (zehanz):

I agree with you that this looks even worse, but from now on it gets better. If you multiply both the numerator and the denominator of the large fraction with \(\cos^2(\frac{\pi}{4}-\theta)\), you will see that the whole thing looks much simpler. Go ahead and try it!

OpenStudy (lynfran):

ok

OpenStudy (zehanz):

After multiplying with \(\cos^2(\frac{\pi}{4}-\theta)\), this is what comes out: \(\dfrac{\cos^2(\frac{\pi}{4}-\theta)-\sin^2(\frac{\pi}{4}-\theta)}{\cos^2(\frac{\pi}{4}-\theta)+\sin^2(\frac{\pi}{4}-\theta)}\)

OpenStudy (zehanz):

What is the next step?

OpenStudy (zehanz):

It simplifies to \(\dfrac{\cos 2(\frac{\pi}{4}-\theta)}{1}=\cos(\frac{\pi}{2}-2\theta)\). Think about what can connect this to the RHS of the identity!

OpenStudy (lynfran):

Sin2theta which =2sin(theta)cos(theta ) thanks alot

OpenStudy (zehanz):

YW!